OpenStudy (anonymous):

if 3 root 5/(root 5 + root 2) = a - root b, find a - root b

OpenStudy (asnaseer):

just rationalise the denominator by multiplying the numerator and denominator by:\[\sqrt{5}-\sqrt{2}\]

OpenStudy (anonymous):

can you solve this question completely

OpenStudy (asnaseer):

do the following multiplication and see what you end up with:\[\frac{3\sqrt{5}}{\sqrt{5}+\sqrt{2}}=\frac{3\sqrt{5}}{\sqrt{5}+\sqrt{2}}\times\frac{\sqrt{5}-\sqrt{2}}{\sqrt{5}-\sqrt{2}}=\frac{3\sqrt{5}(\sqrt{5}-\sqrt{2})}{(\sqrt{5}+\sqrt{2})(\sqrt{5}-\sqrt{2})}\]